# Convexified Message-Passing Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) have become prominent methods for graph representation learning, demonstrating strong empirical results on diverse graph prediction tasks. In this paper, we introduce Convexified Message-Passing Graph Neural Networks (CGNNs), a novel and general framework that combines the power of message-passing GNNs with the tractability of convex optimization. By mapping their nonlinear filters into a reproducing kernel Hilbert space, CGNNs transform training into a convex optimization problem, which can be solved efficiently and optimally by projected gradient methods. This convexity further allows the statistical properties of CGNNs to be analyzed accurately and rigorously. For two-layer CGNNs, we establish rigorous generalization guarantees, showing convergence to the performance of the optimal GNN. To scale to deeper architectures, we adopt a principled layer-wise training strategy. Experiments on benchmark datasets show that CGNNs significantly exceed the performance of leading GNN models, achieving 10–40% higher accuracy in most cases, underscoring their promise as a powerful and principled method with strong theoretical foundations. In rare cases where improvements are not quantitatively substantial, the convex models either slightly exceed or match the baselines, stressing their robustness and wide applicability. Though over-parameterization is often employed to enhance performance in non-convex models, we show that our CGNNs framework yields shallow convex models that can surpass these models in both accuracy and resource efficiency.
@@ 1 Introduction
Graphs serve as versatile models across various domains, such as social networks [47], protein design [22], and medical diagnosis [27]. Graph representation learning has thus attracted widespread attention in recent years, with Graph Neural Networks (GNNs) emerging as powerful tools. Message Passing Neural Networks (MPNNs) [18, 20, 45], a most widely adopted class of GNNs, provide a scalable and elegant architecture that effectively exploits a message-passing mechanism to extract useful information and learn high-quality representations from graph data.
In this paper, we present the new and general framework of Convexified Message-Passing Graph Neural Networks (CGNNs), reducing the training of message-passing GNNs to a convex optimization problem solvable efficiently and optimally via projected gradient methods. The convexity of CGNNs offers a distinct advantage: it ensures global optimality and computational tractability without relying on, e.g., over-parameterization, unlike many existing approaches [1, 10, 11, 60]. This convexity also enables a precise and theoretically grounded analysis of generalization and statistical behavior.
We exhibit our framework by convexifying two-layer Graph Convolution Networks (GCNs) [7, 24], a widely-used GNN model for learning over graph-structured data. Inspired by the convexification of convolutional neural networks by Zhang et al. [59], we adapt their methodology to the graph domain. A high-level overview of our convexification pipeline is summarized in Figure 1, explained as follows. The key challenge lies in handling GCNs’ nonlinearity: we address this by first relaxing the space of GCN filters to a reproducing kernel Hilbert space (RKHS) (step (1) in Figure 1), effectively reducing the model to one with a linear activation function (step (2)). We then show that GCNs with RKHS filters admit a low-rank matrix representation (step (3)). Finally, by relaxing the non-convex rank constraint into a convex nuclear norm constraint (step (4)), we arrive at our CGCN formulation.
@@
Finally, we perform an extensive empirical analysis over various benchmark datasets, where our framework exhibits outstanding results. First, we apply our framework to a wide range of leading message-passing GNN models, including GAT [45], GATv2 [4], GIN [51], GraphSAGE [20], and ResGatedGCN [3]. We also extended our framework to hybrid transformer models, specifically GraphGPS [36] and GraphTrans [40, 49], which integrate local message passing GNNs with global attention mechanisms. Remarkably, in most cases, the convex variants outperform the original non-convex counterparts by a large magnitude, obtaining 10–40% higher accuracy, highlighting the practical power of our approach alongside its theoretical appeal. Even in the minority of cases where improvements are modest, the convexified models still match or slightly exceed the baseline performance, further validating the robustness and broad applicability of our method. While over-parameterization is commonly used to boost performance in non-convex models [1, 10, 11, 60], we show that our CGNNs framework enables the construction of shallow convex architectures that outperform over-parameterized non-convex ones in both accuracy and resource efficiency.
@@ 2 Related Work
Message-Passing GNNs. Early developments of message-passing GNNs include GCN [7, 24], GAT [45], GIN [51], GraphSAGE [20], and others. Such methods effectively leverage a message-passing mechanism to recursively aggregate useful neighbor information [18], enabling them to learn high-quality representations from graph data. Those models have proven effective across diverse tasks, obtaining state-of-the-art performance in various graph prediction tasks (see, e.g., [5, 20, 36, 45, 51, 55]). Our framework of CGNNs harnesses the power of message-passing GNNs by transforming training into a convex optimization problem. This shift not only enables efficient computation of globally optimal solutions, but also allows an accurate and rigorous analysis of statistical properties.
@@
Convergence of Gradient Methods to Global Optima. A vast line of research analyzed the convergence of gradient-based methods to global optima when training neural networks, but under restrictive conditions. Several works rely on over-parameterization, often requiring network widths far exceeding the number of training samples [1, 11, 60], with milder assumptions in shallow networks [10, 43]. However, the convexity of our CGNNs enables global optimality and computational efficiency without relying on any of the above restrictive assumptions.
@@
Convex Neural Networks. The area of convex neural networks is relatively underexplored. Zhang et al. [57] proposed a convex relaxation step for fully-connected neural networks, which inspired the convexification of convolutional neural networks (CNNs) [59]. Other works study linear CNNs [19] and linear GNNs [41], which are of less interest in practice since they collapse all layers into one. Recently, Cohen and Agmon [6] proposed convex versions of specific GNN models termed as aggregation GNN, which apply only to a signal with temporal structure [15]. Their work is narrowly tailored and focused on a specific application in swarm robotics, where convexity arises from modeling assumptions tailored to domain adaptation. In contrast, our CGNN framework establishes a general and theoretically principled convexification of diverse message-passing GNNs by transforming the standard architecture itself, independent of application-specific constraints. Unlike their formulation that does not preserve, e.g., standard GNN modularity, our approach retains these key structural features and applies broadly to a wide range of graph learning tasks. As such, CGNNs represent a scalable and domain-agnostic foundation for convex GNN training, grounded in a fundamentally different mathematical framework.
Graph Transformers. Instead of aggregating local neighborhood, Graph Transformers capture interaction information between any pair of nodes via a single self-attention layer. Some works design specific attention mechanisms or positional encodings [21, 25, 28, 52], while others combine message-passing GNNs to build hybrid architectures [5, 36, 40, 49]. Those methods enable nodes to interact with any other node, facilitating direct modeling of long-range relations. Our framework broadly applies to such hybrid methods, as we exhibit on the leading models GraphGPS [36] and GraphTrans [40, 49]. Empirically, applying our framework consistently improves performance in graph classification tasks, with significant gains on GraphGPS, showcasing the versatility and effectiveness of our approach.
@@ 3 Preliminaries
Network data can be captured by an underlying undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}=\{1,...,n\}$ denotes the vertex set, $\mathcal{E}$ denotes the edge set and $\mathcal{W}:\mathcal{E}\rightarrow\mathbb{R}$ denotes the edge weight function, while a missing edge $(i,j)$ is depicted by $\mathcal{W}(i,j)=0$. The neighborhood of node $i\in\mathcal{V}$ is denoted by $N_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}$. The data on top of this graph is modeled as a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$, whose $i^{\text{th}}$ row $[\mathcal{X}]_{i}=\mathbf{x}_{i}$ is an $F$-dimensional feature vector assigned to node $i\in\mathcal{V}$. We leverage the framework of graph signal processing (GSP) [33] as the mathematical foundation for learning from graph signals. For relating the graph signal $\mathcal{X}$ with the underlying structure of $\mathcal{G}$, we define a graph shift operator (GSO) by $\mathcal{S}\in\mathbb{R}^{n\times n}$, which satisfies $[\mathcal{S}]_{ij}=s_{ij}=0$ if $(j,i)\notin\mathcal{E}$ for $j\neq i$. Commonly utilized GSOs include the adjacency matrix [37, 38], the Laplacian matrix [42], and their normalized counterparts [7, 16]. By viewing $\mathcal{S}$ as a linear operator, the graph signal $\mathcal{X}$ can be shifted over the nodes, yielding that the output at node $i$ for the $f^{\text{th}}$ feature becomes $[\mathcal{S}\mathcal{X}]_{if}=\sum_{j=1}^{n}[\mathcal{S}]_{ij}[\mathcal{X}]_{jf}=\sum_{j\in N_{i}}s_{ij}x_{jf}$. Note that, whereas $\mathcal{S}\mathcal{X}$ corresponds to a local information exchange between a given node and its direct neighbors, applying this linear operator recursively aggregates information from nodes in further hops along the graph. Namely, the $k$-shifted signal $\mathcal{S}^{k}\mathcal{X}$ aggregates information from nodes that are $k$-hops away.
Graph Convolutional Networks (GCNs) [17] stack multiple graph (convolutional) layers to extract high-level node representations from graph signals. To formally describe the creation of convolutional features in a GCN with $L\in\mathbb{N}$ layers, let the output of layer $\ell-1$ (and thus the input to the $\ell^{\text{th}}$ layer) be $\mathcal{X}_{\ell-1}$, comprising of $F_{\ell-1}$ feature vectors. The input to layer 00 is $\mathcal{X}_{0}=\mathcal{X}$, yielding $F_{0}=F$, while the final output has $G:=F_{L}$ features. Given a set of filters $A:=\{\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}|k\in[K]\cup\{0\},\ell\in[L]\}$, the graph (convolutional) filter at layer $\ell$ processes the $F_{\ell-1}$ features of the input graph signal $\mathcal{X}_{\ell-1}$ in parallel by $F_{\ell}$ graph filters to output the following $F_{\ell}$ convolutional features:
@@ 3.1 Training GCNs
We focus on the graph classification task with $G$ classes. Here, we are given a training set $\mathcal{T}:=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$ of $m$ samples, where the $j^{\text{th}}$ sample consists of a graph $\mathcal{G}^{(j)}$, a graph signal $\mathcal{X}^{(j)}\in\mathbb{R}^{n\times F}$ over $\mathcal{G}^{(j)}$ and the graph’s true label $y^{(j)}\in[G]$. Without loss of generality, we assume that each graph $\mathcal{G}^{(j)}$ has $n$ nodes (otherwise, we can add isolated dummy nodes). We then train a GCN (as formulated in (2)), while node features from the last layer are aggregated to obtain the graph-level representation via a permutation invariant function, such as summation or a more sophisticated graph-level pooling function [54, 56]. Training is done by solving the following optimization problem for some loss function $J:\mathbb{R}^{n\times G}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$:

### v2

@@ Abstract
Graph Neural Networks (GNNs) are key tools for graph representation learning, demonstrating strong results across diverse prediction tasks. In this paper, we present Convexified Message-Passing Graph Neural Networks (CGNNs), a novel and general framework that combines the power of message-passing GNNs with the tractability of convex optimization. By mapping their nonlinear filters into a reproducing kernel Hilbert space, CGNNs transform training into a convex optimization problem, which projected gradient methods can solve both efficiently and optimally. Convexity further allows CGNNs’ statistical properties to be analyzed accurately and rigorously. For two-layer CGNNs, we establish rigorous generalization guarantees, showing convergence to the performance of an optimal GNN. To scale to deeper architectures, we adopt a principled layer-wise training strategy. Experiments on benchmark datasets show that CGNNs significantly exceed the performance of leading GNN models, obtaining 10–40% higher accuracy in most cases, underscoring their promise as a powerful and principled method with strong theoretical foundations. In rare cases where improvements are not quantitatively substantial, the convex models either slightly exceed or match the baselines, stressing their robustness and wide applicability. Though over-parameterization is often used to enhance performance in non-convex models, we show that our CGNNs yield shallow convex models that can surpass non-convex ones in accuracy and model compactness.
@@ 1 INTRODUCTION
Graphs serve as versatile models across various domains, such as social networks (Wang et al., 2018), protein design (Ingraham et al., 2019), and medical diagnosis (Li et al., 2020). Graph representation learning has thus attracted widespread attention in recent years, with Graph Neural Networks (GNNs) emerging as powerful tools. Message Passing Neural Networks (MPNNs) (Gilmer et al., 2017; Hamilton et al., 2017; Veličković et al., 2018), a most widely adopted class of GNNs, provide a scalable and elegant architecture that effectively exploits a message-passing mechanism to extract useful information and learn high-quality representations from graph data.
In this paper, we present the new and general framework of Convexified Message-Passing Graph Neural Networks (CGNNs), reducing the training of message-passing GNNs to a convex optimization problem solvable efficiently and optimally via projected gradient methods. The convexity of CGNNs offers a distinct advantage: it ensures global optimality and computational tractability without relying on, e.g., over-parameterization, unlike many existing approaches (Allen-Zhu et al., 2019; Du and Lee, 2018; Du et al., 2018; Zou et al., 2020). This convexity also enables a precise and theoretically grounded analysis of generalization and statistical behavior.
We exhibit our framework by convexifying two-layer Graph Convolution Networks (GCNs) (Defferrard et al., 2016; Kipf and Welling, 2017), a widely-used GNN model for learning over graph-structured data. Inspired by the convexification of convolutional neural networks by Zhang et al. (2017), we adapt their methodology to the graph domain. A high-level overview of our convexification pipeline is summarized in Figure 1, explained as follows. The key challenge lies in handling GCNs’ nonlinearity: we address this by first relaxing the space of GCN filters to a reproducing kernel Hilbert space (RKHS) (step (1) in Figure 1), effectively reducing the model to one with a linear activation function (step (2)). We then show that GCNs with RKHS filters admit a low-rank matrix representation (step (3)). Finally, by relaxing the non-convex rank constraint into a convex nuclear norm constraint (step (4)), we arrive at our CGCN formulation.
@@
Finally, we perform an extensive empirical analysis over various benchmark datasets, where our framework exhibits outstanding results. First, we apply our framework to a wide range of leading message-passing GNN models, including GAT (Veličković et al., 2018), GATv2 (Brody et al., 2022), GIN (Xu et al., 2019), GraphSAGE (Hamilton et al., 2017), and ResGatedGCN (Bresson and Laurent, 2017). We also extended our framework to hybrid transformer models, specifically GraphGPS (Rampášek et al., 2022) and GraphTrans (Shi et al., 2021; Wu et al., 2021), which integrate local message passing GNNs with global attention mechanisms. Remarkably, in most cases, the convex variants outperform the original non-convex counterparts by a large magnitude, obtaining 10–40% higher accuracy, highlighting the practical power of our approach alongside its theoretical appeal. Even in the minority of cases where improvements are modest, the convexified models still match or slightly exceed the baseline performance, further validating the robustness and broad applicability of our method. While over-parameterization is commonly used to boost performance in non-convex models (Allen-Zhu et al., 2019; Du and Lee, 2018; Du et al., 2018; Zou et al., 2020), we show that our CGNNs framework enables the construction of shallow convex architectures that outperform over-parameterized non-convex ones in both accuracy and model compactness.
@@ 2 RELATED WORK
Message-Passing GNNs. Early developments of message-passing GNNs include GCN (Defferrard et al., 2016; Kipf and Welling, 2017), GAT (Veličković et al., 2018), GIN (Xu et al., 2019), and GraphSAGE (Hamilton et al., 2017). Such methods effectively exploit a message-passing mechanism to recursively aggregate useful neighbor information (Gilmer et al., 2017), enabling them to learn high-quality representations from graph data. Those models have proven effective across diverse tasks, obtaining state-of-the-art performance in various graph prediction tasks (see, e.g., (Chen et al., 2022; Hamilton et al., 2017; Rampášek et al., 2022; Veličković et al., 2018; Xu et al., 2019; You et al., 2021)). Our CGNNs harness the power of message-passing GNNs by transforming training into a convex optimization problem. This shift not only enables efficient computation of globally optimal solutions, but also allows an accurate and rigorous analysis of statistical properties. In Appendix A, we provide more details about hybrid transformer models.
@@
Convergence of Gradient Methods to Global Optima. A vast line of research analyzed the convergence of gradient-based methods to global optima when training neural networks, but under restrictive conditions. Several works rely on over-parameterization, often requiring network widths far exceeding the number of training samples (Allen-Zhu et al., 2019; Du et al., 2018; Zou et al., 2020), with milder assumptions in shallow networks (Du and Lee, 2018; Soltanolkotabi et al., 2018). However, the convexity of our CGNNs enables global optimality and computational efficiency without relying on any of the above restrictive assumptions.
@@
Convex Neural Networks. The area of convex neural networks is relatively underexplored. Zhang et al. (2016b) proposed a convex relaxation step for fully-connected neural networks, which inspired the convexification of convolutional neural networks (CNNs) (Zhang et al., 2017); see Appendix A.2 for an extensive comparison to (Zhang et al., 2017). Other works study linear CNNs (Gunasekar et al., 2018) and linear GNNs (Shin et al., 2023), which are of less interest in practice since they collapse all layers into one. Recently, Cohen and Agmon (2021a) proposed convex versions of specific GNN models termed as aggregation GNN, which apply only to a signal with temporal structure (Gama et al., 2018). Their work is narrowly tailored and focused on a specific application in swarm robotics (Cohen and Agmon, 2021c, b, 2020), where convexity arises from modeling assumptions tailored to domain adaptation. In contrast, our CGNN framework establishes a general and theoretically principled convexification of diverse message-passing GNNs by transforming the standard architecture itself, independent of application-specific constraints. Unlike their formulation that does not preserve, e.g., standard GNN modularity, our approach retains these key structural features and applies broadly to a wide range of graph learning tasks. In particular, our contribution resolves the key limitations of Cohen and Agmon (2021a) through three fundamental distinctions: (1) Scope, by enabling training on variable-sized graphs (inductive setting) versus their fixed-topology restriction; (2) Theory, by extending analysis to general non-polynomial filters and proving strictly tighter nuclear norm bounds; and (3) Methodology, by utilizing an RKHS formulation that supports modern optimizers and generic modular tasks. In Appendix A.3, we supply an extensive comparison to Cohen and Agmon (2021a), detailing the fundamental distinctions from their work. As such, CGNNs represent a scalable and domain-agnostic foundation for convex GNN training, grounded in a fundamentally different mathematical framework.
Classical Graph Kernels. Our framework offers a unique theoretical bridge between modern deep graph learning and classical graph kernels. While GNNs have largely superseded kernel methods in popularity, establishing the formal connections between them contextualizes the theoretical advances of the CGNN framework, particularly regarding the role of Reproducing Kernel Hilbert Spaces (RKHS) and optimization landscapes. We supply a detailed comparison to classic graph kernels in Appendix A.5.
GNNs as Unified Optimization Problems. Our framework relates to and fundamentally differs from prior work on GNNs as unified optimization problems. Specifically, Chen et al. (2023b) classify spatial GNNs based on how they aggregate neighbor information, while discussing how graph convolution can be depicted as an optimization problem. They also discuss how filters of spectral GNNs can be approximated via either linear, polynomial or rational approximation. The works by He et al. (2021), Wang and Zhang (2022) and Guo and Wei (2023) focus on polynomial approximation, treating spectral filters as parameterized by a polynomial basis (see Appendix A.4 for details). However, in all those works training remains non-convex. Our contribution is thus orthogonal and complementary: rather than deriving architectures, we convexify the training problem itself for general, non-convex message-passing GNNs (not only spectral ones). Training thus becomes a convex optimization problem, ensuring global optimality as well as an accurate and rigorous analysis of statistical properties. Further, our framework applies directly to general message-passing GNNs and hybrid transformer architectures like GraphGPS and GraphTrans, whereas The works by He et al. (2021), Wang and Zhang (2022) and Guo and Wei (2023) only cover spectral GNNs. In fact, the optimization-oriented analysis of Wang and Zhang (2022) is restricted to linear GNNs, while we focus on convexifying non-convex GNNs. We will add a dedicated discussion explicitly contrasting these settings.
@@ 3 PRELIMINARIES
Network data can be captured by an underlying undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}=\{1,...,n\}$ denotes the vertex set, $\mathcal{E}$ denotes the edge set and $\mathcal{W}:\mathcal{E}\rightarrow\mathbb{R}$ denotes the edge weight function, while a missing edge $(i,j)$ is depicted by $\mathcal{W}(i,j)=0$. The neighborhood of node $i\in\mathcal{V}$ is denoted by $N_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}$. The data on top of this graph is modeled as a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$, whose $i^{\text{th}}$ row $\mathcal{X}]_{i}=\mathbf{x}_{i}$ is an $F$-dimensional feature vector assigned to node $i\in\mathcal{V}$. We leverage the framework of graph signal processing (GSP) (Ortega et al., 2018) as the mathematical foundation for learning from graph signals. For relating the graph signal $\mathcal{X}$ with the underlying structure of $\mathcal{G}$, we define a graph shift operator (GSO) by $\mathcal{S}\in\mathbb{R}^{n\times n}$, which satisfies $\mathcal{S}]_{ij}=s_{ij}=0$ if $(j,i)\notin\mathcal{E}$ for $j\neq i$. Commonly utilized GSOs include the adjacency matrix (Sandryhaila and Moura, 2013, 2014), the Laplacian matrix (Shuman et al., 2013), and their normalized counterparts (Defferrard et al., 2016; Gama et al., 2019a). By viewing $\mathcal{S}$ as a linear operator, the graph signal $\mathcal{X}$ can be shifted over the nodes, yielding that the output at node $i$ for the $f^{\text{th}}$ feature becomes $\mathcal{S}\mathcal{X}]_{if}=\sum_{j=1}^{n}[\mathcal{S}]_{ij}[\mathcal{X}]_{jf}=\sum_{j\in N_{i}}s_{ij}x_{jf}$. Note that, whereas $\mathcal{S}\mathcal{X}$ corresponds to a local information exchange between a given node and its direct neighbors, applying this linear operator recursively aggregates information from nodes in further hops along the graph. Namely, the $k$-shifted signal $\mathcal{S}^{k}\mathcal{X}$ aggregates information from nodes that are $k$-hops away.
Graph Convolutional Networks (GCNs) stack multiple graph (convolutional) layers to extract high-level node representations from graph signals (Gama et al., 2019b). To formally describe the creation of convolutional features in a GCN with $L\in\mathbb{N}$ layers, let the output of layer $\ell-1$ (and thus the input to the $\ell^{\text{th}}$ layer) be $\mathcal{X}_{\ell-1}$, comprising of $F_{\ell-1}$ feature vectors. The input to layer 0 is $\mathcal{X}_{0}=\mathcal{X}$, yielding $F_{0}=F$, while the final output has $G:=F_{L}$ features. Given a set of filters $A:=\{\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}|k\in[K]\cup\{0\},\ell\in[L]\}$, the graph (convolutional) filter at layer $\ell$ processes the $F_{\ell-1}$ features of the input graph signal $\mathcal{X}_{\ell-1}$ in parallel by $F_{\ell}$ graph filters to output the following $F_{\ell}$ convolutional features:
@@ 3.1 Training GCNs
We focus on the graph classification task with $G$ classes. Here, we are given a training set $\mathcal{T}:=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$ of $m$ samples, where the $j^{\text{th}}$ sample consists of a graph $\mathcal{G}^{(j)}$, a graph signal $\mathcal{X}^{(j)}\in\mathbb{R}^{n\times F}$ over $\mathcal{G}^{(j)}$ and the graph’s true label $y^{(j)}\in[G]$. Without loss of generality, we assume that each graph $\mathcal{G}^{(j)}$ has $n$ nodes (otherwise, we can add isolated dummy nodes). We then train a GCN (as formulated in (2)), while node features from the last layer are aggregated to obtain the graph-level representation via a fixed permutation invariant function, such as summation or a more sophisticated graph-level pooling function (Ying et al., 2018b; Zhang et al., 2018). Training is done by solving the following optimization problem for some loss function $J:\mathbb{R}^{n\times G}\times[G]\rightarrow\mathbb{R}$:
@@
Remark 3.1.
As our focus is on understanding generalization of convexified message-passing GNNs, we restricted evaluations to graph classification. This is exactly as in prior works that only consider a single task (see, e.g., (Lachi et al., 2025; Chen et al., 2023a; Wu et al., 2022)). However, our convexification framework directly extends to node-level prediction, by using the node features from the last layer instead of aggregating them to obtain the graph-level representation. While this results in a node-level instead of graph-level loss, our convexification procedure remains identical.